### v3

# Conjugacy Class Growth in Virtually Abelian Groups

Abstract.
We study the growth of conjugacy classes in finitely generated virtually abelian groups. That is, the number of elements in the ball of radius $n$ in the Cayley graph which intersect a fixed conjugacy class. In the class of virtually abelian groups, we prove that this function is always asymptotically equivalent to a polynomial. Furthermore, we show that in any affine Coxeter group, the degree of polynomial growth of a conjugacy class is equivalent to the reflection length of any element of that class.
2020 Mathematics Subject Classification. 05E16, 20E45, 20F55, 20F69.
1. Introduction
Given a finitely generated group $G$ generated by some set $S$, every element in $G$ can be represented by some word over the alphabet $S\cup S^{-1}$. For an element $g\in G$, we define the length $\ell_{S}(g)$ of $g$ to be the minimal length of all words which represent $g$ with respect to the generating set $S$. This naturally leads to the study of the growth function of $G$, which counts the number of elements of length at most $n$. The asymptotic properties of growth functions of groups have been extensively studied with a highlight being Gromov’s celebrated theorem [Gromov] characterising groups of polynomial growth as the virtually nilpotent groups.
The growth function can be generalised by restricting our attention to subsets of interest. In particular, taking some subset $U$ of $G$, we let $\beta_{U,S}(n)$ be the number of elements in $U$ whose length is at most $n$ and look at how this number changes as $n$ increases. The growth of subgroups of groups has been studied by a number of authors [CantrellSharp, DavisOlshanskii, Schesler] but much less attention has been paid to other interesting subsets.
In this article we focus our attention on conjugacy classes in virtually abelian groups. In particular, we take a conjugacy class $C$ and we look at the growth function of $C$ as a subset of $G$. To this end, we prove the following theorem.
Theorem LABEL:thm:conj_class_poly.
Let $G$ be a finitely generated virtually abelian group. Then every conjugacy class has polynomial growth.
We emphasise that this is not to be confused with the conjugacy growth function, which counts the number of conjugacy classes intersecting the ball of radius $n$. The conjugacy growth function in virtually abelian groups was studied in [Evetts]. Combinatorial properties including growth of other subsets of virtually abelian groups are explored in [RationalVAb], and as noted in Remark LABEL:rem:altproof, Theorem LABEL:thm:conj_class_poly also follows from general results in that article. However, the method of proof we use here is much more direct, and as we will see, allows for the degree of polynomial growth to be explicitly calculated.
To the authors’ knowledge, the first appearance of the growth of a conjugacy class is in [ParkPaul], where it appears as a special case of the ‘orbital counting problem in conjugacy classes’. The authors of that paper are interested in examples of groups where all non-trivial conjugacy classes have equivalent growth functions, as well as the exponential growth rate of the function. They show that in the cases of hyperbolic groups, and the integer Heisenberg group, all infinite conjugacy classes have the same growth. In the present paper we show that this is not the case in general for virtually abelian groups (see Proposition LABEL:prop:allgrowths), while providing an example of where it does hold (Example LABEL:ex:kleinbottle).
As an example of the growth function of conjugacy classes in virtually abelian groups, we consider the affine Coxeter groups. It turns out that the polynomial growth is directly related to the reflection length. For an element $w$ in an affine Coxeter group $W$ with simple reflections $S$, we decompose $w$ into a translation part $t$ and a finite part $u$ where $w=tu$. The reflection length $\ell_{R}(w)$ of $w$ is the length with regards to the set of all reflections instead of just the simple reflections. This leads us to our second main theorem of this article.
Theorem LABEL:thm:poly_growth_Coxeter.
Let $W$ be an affine Coxeter group and $w=tu$ an arbitrary element in $W$ decomposed as a translation part and a finite part. Then the growth of the conjugacy class with respect to the generating set $S$ is equivalent to $n^{\ell_{R}(u)}$.
This article is organised in the following way. We start by giving preliminary definitions and basic results in group theory and growth in Section 2. In Section LABEL:sec:coxeter_groups, we give background definitions and results on affine Coxeter groups. The affine Coxeter groups are a particularly nice class of virtually abelian groups as they have a rich geometric structure to them. Section LABEL:sec:virtually_abelian_groups is then devoted to showing our two main results: (1) that the growth function of conjugacy classes in virtually abelian groups is polynomial and (2) that the growth function of the conjugacy class of an element in an affine Coxeter group is polynomial with degree equal to the reflection length of its finite part.
2. Definitions and basic results
We begin by recalling some standard definitions in group theory and growth functions. The interested reader is referred to [Loh] and [Mann] for a more thorough introduction to the growth of groups.
2.1. Growth in groups
We start by giving some basic notions on words representing elements of a group and the growth rate of a set in relation to a group. {defi} Let $G$ be a finitely generated group and $S$ a choice of finite generating set. We let $\left\{S\cup S^{-1}\right\}^{\star}$ denote the words over the alphabet $S$ and its inverses, including the empty word. Given a word $w\in\left\{S\cup S^{-1}\right\}^{\star}$ we let $\lvert w\rvert$ denote the length of the word $w$ and we write $w=_{G}g$ if $w$ represents $g$ in the group $G$. Many words can represent the same element, but there will always be some word of minimal length representing a given $g$. The word length of an element $g\in G$ with respect to $S$ is then defined to be this minimal length
$\ell_{S}(g)=\min\left\{\lvert w\rvert\;\colon\;w\in\{S\cup S^{-1}\}^{*},w=_{G}g\right\}.$
If there is no ambiguity in the generating set, we will write $\ell(g)$ as a shorthand for $\ell_{S}(g)$. The length of elements can be extended to the word metric on $G$ via the distance function given by $d_{G,S}(g,h)=\ell_{S}(g^{-1}h)$. We write $B_{G,S}(n)=\left\{g\in G\;\colon\;\ell_{S}(g)\leq n\right\}$ for the metric $n$-ball in $G$ with respect to $S$. If there is no ambiguity in the group or generating set, we will use $B(n)$ to represent $B_{G,S}(n)$. Let $U$ be any subset of $G$. We define the growth of $U$ in $G$ as follows:
$\beta_{U,S}(n)\coloneqq\lvert\left\{g\in U\;\colon\;\ell_{S}(g)\leq n\right\}\rvert=\lvert U\cap B_{G,S}(n)\rvert.$
As before, if there is no ambiguity in the set $S$, the group $G$ or the subset $U$, instead of $\beta_{U,S}(n)$ we will write $\beta_{U}(n)$. When $U=G$ this is the familiar notion of the (cumulative) growth function of a group.
The following lemma shows that the number of elements in the $n$-ball is the same, regardless of the choice of centre point.
Lemma 1.
Fix any element $h\in G$. Then |$\lvert\left\{g\in G\;\colon\;d_{G,S}(h,g)\leq n\right\}\rvert=\beta_{G}(n)$.
Proof 2.1.
For $h,x,y\in G$ we have $d_{G,S}(x,y)=\ell_{S}(x^{-1}y)=\ell_{S}(x^{-1}h^{-1}hy)=d_{G,S}(hx,hy)$, and so the word metric is invariant under left multiplication. In other words $d_{G,S}(1,x)\leq n$ if and only if $d_{G,S}(h,hx)\leq n$ and we have a length-preserving bijection between the elements of the $n$-ball centred at $1$ and the elements of the $n$-ball centred at $h$.
Note that the growth of a subgroup $U$ of $G$ is often referred to as relative growth, since the subgroup may itself be finitely generated and therefore have an intrinsic notion of growth. In this paper, we study subsets that are not necessarily subgroups and so we drop the term ‘relative’ since the natural growth is the one inherited from the ambient group.
We recall the usual notion of equivalence of growth functions. Let $f,g\colon{\mathbb{N}}\rightarrow{\mathbb{N}}$ be two functions. We write $f\preccurlyeq g$ if there exists $\lambda\geq 1$ such that
$f(n)\leq\lambda g(\lambda n+\lambda)+\lambda$
for all $n\in{\mathbb{N}}$. If $f\preccurlyeq g$ and $g\preccurlyeq f$ then we write $f\sim g$ and say that the functions are
Definition 2.
equivalent. Note that this defines an equivalence relation. We write $f\prec g$ if $f\preccurlyeq g$ but it is not the case that $f\sim g$.
It is a standard fact that, up to equivalence, the growth of a group does not depend on the choice of finite generating set (since every element written as a word of one generating set can be written as a word of bounded length in another generating set). The following lemma is easily proved in the same way.
Lemma 3.
Up to equivalence, the growth of a subset $U$ in $G$ does not depend on the choice of generating set for $G$.
We say that a subset $U$ has